OpenStudy (17bohtal):

The length of a rectangle is 3 centimeters less than twice its width. The perimeter of he rectangles is 48 cm. What is the dimensions of the rectangle?

OpenStudy (anonymous):

Let the length of the rectangle be l and width be w Since it is given that the length is 3cm less than twice the width we have l=2w-3 But Perimeter P=2(w+l) 48=2(w+l) Dividing both sides by 2 we get 24=l+w 28=2w-3+w solving this we will get the value of w. Substituting in l=2w-3 we can find l Does this help :)

OpenStudy (anonymous):

did u get it or u need more?